OpenStudy (anonymous):

If I buy 2 apples and 6 oranges, I will spend $5.60. However, if I buy 6 apples and 2 oranges, I will spend $4.80. How much does one orange cost? A. $0.55 B. $0.70 C. $0.75 D. $0.60

hero (hero):

2a + 6o = 5.60 6a + 2o =4.80 Write both equations in terms of a Set a = a Solve for o Solve for a

OpenStudy (anonymous):

i don't understand it that way , im sorry . and im in a rush to get everything done .

hero (hero):

a = (5.60 - 6o)/2 a = (4.80 - 2o)/6 a = a 6(5.60 - 6o) = 2(4.80 - 2o) 33.60 - 36o = 9.60 - 4o 24 = 32o o = .75 a = .55